A heavy ball is suspended from the ceiling of a motor car through a light string.A transverse pulse Travels at a speed of 60 centimetre per second on the string when the car is at rest and 62 centimetre per second when the car accelerates on a horizontal Road find the acceleration of the car.

Grade:12th pass

1 Answers

pranav gupta
6 years ago
When the car is at rest :T = mg V = (T/u)^0.5(60/100)^2 = mg/(m/l)L = 3.6 cm When car is accelerating :T^2 = (mg)^2 + (ma)^2Now V = (T/u)^0.5V is 62/100 and u is m/l Solving we get :a = 3.7 m/s^2